# Goal-oriented time adaptivity for port-Hamiltonian systems

### v1

@@ Abstract.
Port-Hamiltonian systems provide an energy-based modeling paradigm for dynamical input-state-output systems. At their core, they fulfill an energy balance relating stored, dissipated and supplied energy. To accurately resolve this energy balance in time discretizations, we propose an adaptive grid refinement technique based on a posteriori error estimation. The evaluation of the error estimator includes the computation of adjoint sensitivities. To interpret this adjoint equation as a backwards-in-time equation, we show piecewise weak differentiability of the dual variable. Then, leveraging dissipativity of the port-Hamiltonian dynamics, we present a parallelizable approximation of the underlying adjoint system in the spirit of a block-Jacobi method to efficiently compute error indicators. We illustrate the performance of the proposed scheme by means of numerical experiments showing that it yields a smaller violation of the energy balance when compared to uniform refinements and traditional step-size controlled time stepping.
@@ 2. Goal-oriented error estimation for port-Hamiltonian systems
for all $s,t\in[0,T]$, where $R^{1/2}$ denotes the Cholesky factor of $R=R^{\top}\geq 0$.
@@ 2.1. Refinements targeting the energy balance
$I_{\mathrm{glob}}(\tilde{x}):=\left|\int_{0}^{T}\!-u(t)^{\top}\tilde{y}(t)+\tilde{x}(t)^{\top}RQR\tilde{x}(t)\,\mathrm{d}t\;+\;H(\tilde{x}(T))-H(\tilde{x}(0))\right|^{2}.$
@@
$I^{i}_{\mathrm{loc}}(\tilde{x}):=\left|\int_{t_{i}}^{t_{i+1}}\!\!-u(t)^{\top}\tilde{y}(t)+\tilde{x}(t)^{\top}RQR\tilde{x}(t)\,\mathrm{d}t\;+\;H(\tilde{x}(t_{i+1}))-H(\tilde{x}(t_{i}))\right|^{2}$ (2.6)
@@
To ensure that not only the energy balance is satisfied, but to also control the error of the solution in norm, we further consider a weighted quantity of interest, i.e., for $\rho\geq 0$, we set
@@ 2.2. Fundamentals of goal-oriented adapativity
$I(x)-I(x_{h})\approx\langle\mathcal{A}(x-x_{h}),\lambda\rangle_{X^{*},X}=\langle\mathcal{A}(x-x_{h}),\lambda-\lambda_{h}\rangle_{X^{*},X}$ (2.13)
@@ 3.1. State equation
$H^{1}_{\mathcal{T}}([0,T],\mathbb{R}^{n}):=\{v\in L^{2}([0,T],\mathbb{R}^{n})\,\bigl{|}\,v|_{\mathbb{I}_{i}}\in H^{1}([t_{i-1},t_{i}],\mathbb{R}^{n})\ \forall 1\leq i\leq M,\ v(0)\in\mathbb{R}^{n}\}.$
@@
$\lambda^{-}(t_{i}):=\lim_{t\to 0^{+}}\lambda(t_{i}-t)\quad\text{(for }\,0<i\leq M),\qquad\lambda^{+}(t_{i}):=\lim_{t\to 0^{+}}\lambda(t_{i}+t)\quad\text{(for }\,0\leq i<M)$
@@
$E:H^{1}([0,T],\mathbb{R}^{n})$ $\to L^{2}([0,T],\mathbb{R}^{n})\times\mathbb{R}^{n}\times\ldots,\times\mathbb{R}^{n}$
@@
$\langle A^{*}(\lambda,\lambda(0)),x\rangle_{H^{1}_{\mathcal{T}}([0,T],\mathbb{R}^{n})^{*},H^{1}_{\mathcal{T}}([0,T],\mathbb{R}^{n})}=\langle\mathcal{A}^{-}\lambda,(x,x(0),\ldots,x(T))\rangle_{L^{2}(0,T;\mathbb{R}^{n})\times\mathbb{R}^{n}\times\ldots\times\mathbb{R}^{n}}.$
@@ 3.3. Discretization
$\langle\mathcal{A}x,(\varphi,\varphi(0))\rangle_{L^{2}([0,T],\mathbb{R}^{n})\times\mathbb{R}^{n}}$ $=\int_{0}^{T}\langle\dot{x}(t)-Ax(t),\varphi(t)\rangle\,\mathrm{d}t+\langle x(0),\varphi(0)\rangle$
@@
$X_{h}:=\big{\{}v_{h}\in L^{2}([0,T],\mathbb{R}^{n})\,|\,v_{h}|_{\mathbb{I}_{i}}=v_{h,i}\in\mathbb{R}^{n},\quad 1\leq i\leq M\ \mathrm{and}\ v_{h}(0)\in\mathbb{R}^{n}\big{\}}.$ (3.5)
@@
$x_{h}^{-}(t_{i})=\lim_{t\to 0^{+}}x_{h}(t_{i}-t)=x(t_{i}),\qquad x_{h}^{+}(t_{i})=\lim_{t\to 0^{+}}x_{h}(t_{i}+t)=x(t_{i+1})$
@@
$[x_{h}]_{i}:=x_{h}^{+}(t_{i})-x_{h}^{-}(t_{i})=x(t_{i+1})-x(t_{i}).$
@@
$\quad=-\sum_{i=1}^{M}\int_{\mathbb{I}_{i}}\langle x_{h}(t),\dot{\varphi}(t)\rangle\,\mathrm{d}t+\langle x_{h}(T),\varphi(T)\rangle$
$\quad=\sum_{i=1}^{M}\int_{\mathbb{I}_{i}}\langle\dot{x}_{h}(t),\varphi(t)\rangle\,\mathrm{d}t+\sum_{i=1}^{M}\left(-\langle x_{h}^{-}(t_{i}),\varphi(t_{i})\rangle+\langle x_{h}^{+}(t_{i-1}),\varphi(t_{i-1})\rangle\right)+\langle x_{h}(T),\varphi(T)\rangle$
$\quad=\sum_{i=1}^{M-1}\langle x_{h}^{+}(t_{i}),\varphi(t_{i})\rangle-\langle x_{h}^{-}(t_{i}),\varphi(t_{i})\rangle+\langle x_{h}^{+}(0),\varphi^{+}(0)\rangle$
$\quad=\sum_{i=1}^{M-1}\langle x_{h}^{+}(t_{i})-x_{h}^{-}(t_{i}),\varphi(t_{i})\rangle+\langle x_{h}^{+}(0),\varphi(0)\rangle$
$\quad=\sum_{i=0}^{M-1}\langle x_{h}^{+}(t_{i})-x_{h}^{-}(t_{i}),\varphi(t_{i})\rangle+\langle x_{h}^{-}(0),\varphi(0)\rangle$
$\quad=\sum_{i=0}^{M-1}\langle[x_{h}]_{i},\varphi(t_{i})\rangle+\langle x_{h}^{-}(0),\varphi(0)\rangle.$
@@
$\langle\mathcal{A}_{h}x_{h},\varphi_{h}\rangle_{X_{h}^{*},X_{h}}:=\sum_{i=1}^{M}h_{i}\langle-Ax_{h}(t_{i}),\varphi_{h}(t_{i})\rangle+\sum_{i=0}^{M-1}\langle[x_{h}]_{i},\varphi_{h}(t_{i+1})\rangle+\langle x_{h}(0),\varphi_{h}(0)\rangle.$
@@
Note that, due to dissipativity of $A=(J-R)Q$, the diagonal blocks of the matrix on the left hand side are invertible for any choice of stepsizes.
@@
$\langle\mathcal{A}_{h}^{-}\lambda_{h},\varphi_{h}\rangle_{X_{h}^{*},X_{h}}:=\sum_{i=1}^{M}h_{i}\langle-A^{\top}\lambda_{h}(t_{i}),\varphi_{h}(t_{i})\rangle+\sum_{i=0}^{M-1}\langle-[\lambda_{h}]_{i},\varphi_{h}(t_{i})\rangle+\langle\lambda_{h}(T),\varphi_{h}(T)\rangle.$ (3.7)
@@ 4. Implementation of goal-oriented adaptivity
Having defined a variational framework for the formulation of state and adjoint equation, we provide implementation details of the goal-oriented scheme sketched in Section 2. In the following, we consider the state $x\in X:=H^{1}([0,T],\mathbb{R}^{n})$ solving the continuous problem (2.1a), i.e., in view of Proposition 1,
$\langle\mathcal{A}x,(\varphi,\varphi(0))\rangle_{L^{2}([0,T],\mathbb{R}^{n})\times\mathbb{R}^{n}}=\int_{0}^{T}\langle Bu(t),\varphi(t)\rangle\,\mathrm{d}t+\langle x^{0},\varphi(0)\rangle\qquad\forall\varphi\in H^{1}([0,T],\mathbb{R}^{n}).$
@@
$\qquad\qquad\qquad=\langle I_{\mathrm{loc}}^{\prime}(x_{h}),\varphi\rangle_{H^{1}([0,T],\mathbb{R}^{n})^{*},H^{1}([0,T],\mathbb{R}^{n})}\qquad\forall\varphi\in H^{1}([0,T],\mathbb{R}^{n}).$ (4.1)
@@
$\langle I^{\prime}_{\mathrm{loc}}(x_{h}),\varphi\rangle_{X^{*},X}=\sum_{i=0}^{M-1}\bigg{(}\int_{t_{i}}^{t_{i+1}}-u(t)^{\top}y(t)+x(t)^{\top}RQRx(t)\,\mathrm{d}t+H(x(t_{i+1}))-H(x(t_{i}))\bigg{)}\times$
$\qquad\qquad\bigg{(}\int_{t_{i}}^{t_{i+1}}u(t)B^{\top}Q\varphi(t)+2x(t)^{\top}RQR\varphi(t))\,\mathrm{d}t+H^{\prime}(x(t_{i+1}))\varphi(t_{i+1})-H^{\prime}(x(t_{i}))\varphi(t_{i})\bigg{)},$
@@
$\langle\mathcal{A}_{h}^{-}\lambda,\varphi_{h}\rangle_{X_{h}^{*},X_{h}}=\langle I^{\prime}(x_{h}),\varphi_{h}\rangle_{X_{h}^{*},X_{h}}\qquad\forall\varphi_{h}\in X_{h}$ (4.2)
@@
$\left(\begin{smallmatrix}I-h_{1}A^{\top}&-I&0&0&\cdots&0\\ 0&I-h_{2}A^{\top}&-I&0&\cdots&0\\ 0&0&I-h_{3}A^{\top}&-I&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\vdots&\vdots&0&I-h_{M}A^{\top}&-I\\ 0&0&0&0&0&I\\ \end{smallmatrix}\right)\begin{pmatrix}\lambda_{0}\\ \lambda_{1}\\ \vdots\\ \lambda_{M}\end{pmatrix}=\begin{pmatrix}\langle I^{\prime}_{\mathrm{loc}}(x),\varphi_{1}\rangle_{X_{h}^{*},X_{h}}\\ \langle I^{\prime}_{\mathrm{loc}}(x),\varphi_{2}\rangle_{X_{h}^{*},X_{h}}\\ \vdots\\ \langle I^{\prime}_{\mathrm{loc}}(x),\varphi_{M}\rangle_{X_{h}^{*},X_{h}}\end{pmatrix}.$ (4.3)
@@
$\langle I^{\prime}_{\mathrm{loc}}(x),\varphi_{h}\rangle_{X_{h}^{*},X_{h}}\approx\sum_{i=1}^{M-1}$ $\bigg{(}h_{i}(-u_{i}^{\top}y_{i}+x_{i}^{\top}RQRx_{i})+H(x_{i+1})-H(x_{i})\bigg{)}\times$
$(h_{i}(u_{i}B^{\top}Q\varphi_{i}+2x_{i}^{\top}RQR\varphi_{i})+H^{\prime}(x_{i+1})\varphi_{i+1}-H^{\prime}(x_{i})\varphi_{i}).$
@@
$(I_{h}^{1}\lambda-\lambda)(t_{i})=\lambda_{h}(t_{i})-\lambda_{h}(t_{i-1}).$
@@
$\qquad=\sum_{i=1}^{M}\frac{h_{i}}{2}\langle-Ax_{h}(t_{i}),\lambda_{h}(t_{i})-\lambda_{h}(t_{i-1})\rangle+\langle x_{h}(t_{i})-x_{h}(t_{i-1}),\lambda_{h}(t_{i})-\lambda_{h}(t_{i-1})\rangle$
$\qquad\qquad+\frac{h_{i}}{2}\left(\lambda(t_{i-1})^{\top}Bu(t_{i-1})-\lambda(t_{i})^{\top}Bu(t_{i})\right).$
@@
$\begin{split}\eta_{i}&:=\Big{|}\frac{h_{i}}{2}\langle-Ax_{h}(t_{i}),\lambda_{h}(t_{i})-\lambda_{h}(t_{i-1})\rangle+\langle x_{h}(t_{i})-x_{h}(t_{i-1}),\lambda_{h}(t_{i})-\lambda_{h}(t_{i-1})\rangle\\ &\qquad+\frac{h_{i}}{2}\left(\lambda(t_{i-1})^{\top}Bu(t_{i-1})-\lambda(t_{i})^{\top}Bu(t_{i})\right)\Big{|}.\end{split}$ (4.4)
@@
There, we observe that the main additional computational effort compared to usual time stepping is the computation of the adjoint state. Thus, in the following section, we provide a parallelizable approximation for the the backward equation (4.3).

### v2

@@ Abstract.
Port-Hamiltonian systems provide an energy-based modeling paradigm for dynamical input-state-output systems. At their core, they fulfill an energy balance relating stored, dissipated and supplied energy. To accurately resolve this energy balance in time discretizations, we propose an adaptive grid refinement technique based on a posteriori error estimation. The evaluation of the error estimator includes the computation of adjoint sensitivities. To interpret this adjoint equation as a backwards-in-time equation, we show piecewise weak differentiability of the dual variable. Then, leveraging dissipativity of the port-Hamiltonian dynamics, we present a parallelizable approximation of the underlying adjoint system in the spirit of a block-Jacobi method to efficiently compute error indicators. We illustrate the performance of the proposed scheme by means of numerical experiments showing that it yields a smaller violation of the energy balance when compared to uniform refinements and traditional step size controlled time stepping.
This work was supported within the CRC 1701 ”Port-Hamiltonian Systems”.
@@ 2. Goal-oriented error estimation for port-Hamiltonian systems
for all $s,t\in[0,T]$, where $R^{1/2}$ denotes the matrix square root of $R=R^{\top}\geq 0$.
@@ 2.1. Refinements targeting the energy balance
$I_{\mathrm{glob}}(\tilde{x}):=\left|\int_{0}^{T}\!-u(t)^{\top}\tilde{y}(t)+\tilde{x}(t)^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}QRQ}\tilde{x}(t)\,\mathrm{d}t\;+\;H(\tilde{x}(T))-H(\tilde{x}(0))\right|^{2}.$
@@
$I^{i}_{\mathrm{loc}}(\tilde{x}):=\left|\int_{t_{i}}^{t_{i+1}}-u(t)^{\top}\tilde{y}(t)+\tilde{x}(t)^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}QRQ}\tilde{x}(t)\,\mathrm{d}t+H(\tilde{x}(t_{i+1}))-H(\tilde{x}(t_{i}))\right|^{2}$ (2.6)
@@
To ensure that not only the energy balance is satisfied, but to also guarantee a small error in norm, we further consider a weighted quantity of interest, i.e., for $\rho\geq 0$, we set
@@
In view of (2.5), this QoI additionally yields error control in view of the norm.
@@ 2.2. Fundamentals of goal-oriented adapativity
$I(x)-I(x_{h})\approx\langle\mathcal{A}(x-x_{h}),\lambda\rangle_{X^{*},X}=\langle\mathcal{A}(x-x_{h}),\lambda-\lambda_{h}\rangle_{X^{*},X},$ (2.13)
@@ 3.1. State equation
$H^{1}_{\mathcal{T}}([0,T],\mathbb{R}^{n})\!:=\!\{v\in L^{2}([0,T],\mathbb{R}^{n})\,\bigl{|}\,v|_{\mathbb{I}_{i}}\in H^{1}(\mathbb{I}_{i},\mathbb{R}^{n})\ \forall 1\leq i\leq M,v(0)\in\mathbb{R}^{n}\}.$
@@
$\lambda^{-}(t_{i})$ $:=\lim_{t\to 0^{+}}\lambda(t_{i}-t)\quad\text{(for }\,0<i\leq M),$
$\lambda^{+}(t_{i})$ $:=\lim_{t\to 0^{+}}\lambda(t_{i}+t)\quad\text{(for }\,0\leq i<M)$
@@
$E:H^{1}([0,T],\mathbb{R}^{n})$ $\to L^{2}([0,T],\mathbb{R}^{n})\times\mathbb{R}^{n}\times\ldots\times\mathbb{R}^{n}$
@@
$\langle A^{*}(\lambda,\lambda(0)),x$ $\rangle_{H^{1}_{\mathcal{T}}([0,T],\mathbb{R}^{n})^{*},H^{1}_{\mathcal{T}}([0,T],\mathbb{R}^{n})}$
$=\langle\mathcal{A}^{-}\lambda,(x,x(0),\ldots,x(T))\rangle_{L^{2}(0,T;\mathbb{R}^{n})\times\mathbb{R}^{n}\times\ldots\times\mathbb{R}^{n}}.$
@@ 3.3. Discretization
$\langle\mathcal{A}x,$ $(\varphi,\varphi(0))\rangle_{L^{2}([0,T],\mathbb{R}^{n})\times\mathbb{R}^{n}}$
$=\int_{0}^{T}\langle\dot{x}(t)-Ax(t),\varphi(t)\rangle\,\mathrm{d}t+\langle x(0),\varphi(0)\rangle$
@@
$X_{h}\!:=\!\big{\{}v_{h}\in L^{2}([0,T],\mathbb{R}^{n})\,|\,v_{h}|_{\mathbb{I}_{i}}=v_{h,i}\in\mathbb{R}^{n},1\leq i\leq M,\ v_{h}(0)\in\mathbb{R}^{n}\big{\}}.$ (3.5)
@@
$x_{h}^{-}(t_{i})=\lim_{t\to 0^{+}}x_{h}(t_{i}-t)={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{h}}(t_{i}),\qquad x_{h}^{+}(t_{i})=\lim_{t\to 0^{+}}x_{h}(t_{i}+t)={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{h}}(t_{i+1})$
@@
$[x_{h}]_{i}:=x_{h}^{+}(t_{i})-x_{h}^{-}(t_{i})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{h}}(t_{i+1})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{h}}(t_{i}).$
@@
$\qquad=-\sum_{i=1}^{M}\int_{\mathbb{I}_{i}}\langle x_{h}(t),\dot{\varphi}(t)\rangle\,\mathrm{d}t+\langle x_{h}(T),\varphi(T)\rangle$
$\qquad=\sum_{i=1}^{M}\int_{\mathbb{I}_{i}}\langle\dot{x}_{h}(t),\varphi(t)\rangle\,\mathrm{d}t+\sum_{i=1}^{M}\left(-\langle x_{h}^{-}(t_{i}),\varphi(t_{i})\rangle+\langle x_{h}^{+}(t_{i-1}),\varphi(t_{i-1})\rangle\right)$
$\hskip 56.9055pt+\langle x_{h}(T),\varphi(T)\rangle$
$\qquad=\sum_{i=1}^{M-1}\langle x_{h}^{+}(t_{i}),\varphi(t_{i})\rangle-\langle x_{h}^{-}(t_{i}),\varphi(t_{i})\rangle+\langle x_{h}^{+}(0),\varphi^{+}(0)\rangle$
$\qquad=\sum_{i=1}^{M-1}\langle x_{h}^{+}(t_{i})-x_{h}^{-}(t_{i}),\varphi(t_{i})\rangle+\langle x_{h}^{+}(0),\varphi(0)\rangle$
$\qquad=\sum_{i=0}^{M-1}\langle x_{h}^{+}(t_{i})-x_{h}^{-}(t_{i}),\varphi(t_{i})\rangle+\langle x_{h}^{-}(0),\varphi(0)\rangle$
$\qquad=\sum_{i=0}^{M-1}\langle[x_{h}]_{i},\varphi(t_{i})\rangle+\langle x_{h}^{-}(0),\varphi(0)\rangle.$
@@
$\langle\mathcal{A}_{h}x_{h},\varphi_{h}\rangle_{X_{h}^{*},X_{h}}:=\sum_{i=1}^{M}h_{i}\langle-Ax_{h}(t_{i}),\varphi_{h}(t_{i})\rangle$ $+\sum_{i=0}^{M-1}\langle[x_{h}]_{i},\varphi_{h}(t_{i+1})\rangle$
$+\langle x_{h}(0),\varphi_{h}(0)\rangle.$
@@
We emphasize that due to dissipativity of $A=(J-R)Q$, the diagonal blocks of the matrix on the left hand side are invertible for any choice of step sizes.
@@
$\begin{split}\langle\mathcal{A}_{h}^{-}\lambda_{h},\varphi_{h}\rangle_{X_{h}^{*},X_{h}}:=\sum_{i=1}^{M}h_{i}\langle-A^{\top}\lambda_{h}(t_{i}),\varphi_{h}(t_{i})\rangle&+\sum_{i=0}^{M-1}\langle-[\lambda_{h}]_{i},\varphi_{h}(t_{i})\rangle\\ &+\langle\lambda_{h}(T),\varphi_{h}(T)\rangle\end{split}$ (3.7)
@@ 4. Implementation of goal-oriented adaptivity
Having defined a variational framework for the formulation of state and adjoint equation, we provide implementation details of the goal-oriented scheme sketched in Section 2. In the following, we consider the state $x\in X:=H^{1}([0,T],\mathbb{R}^{n})$ solving the continuous problem (2.1a), i.e., in view of Proposition 1, for all $\varphi\in H^{1}([0,T],\mathbb{R}^{n})$
$\langle\mathcal{A}x,(\varphi,\varphi(0))\rangle_{L^{2}([0,T],\mathbb{R}^{n})\times\mathbb{R}^{n}}=\int_{0}^{T}\langle Bu(t),\varphi(t)\rangle\,\mathrm{d}t+\langle x^{0},\varphi(0)\rangle.$
@@
$\qquad\qquad=\langle I_{\mathrm{loc}}^{\prime}(x_{h}),\varphi\rangle_{H^{1}([0,T],\mathbb{R}^{n})^{*},H^{1}([0,T],\mathbb{R}^{n})}\quad\forall\varphi\in H^{1}([0,T],\mathbb{R}^{n}).$ (4.1)
@@
$\langle I^{\prime}_{\mathrm{loc}}(x_{h}),\varphi\rangle_{X^{*},X}$
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}\sum_{i=0}^{M-1}\bigg{(}\int_{t_{i}}^{t_{i+1}}-u(t)^{\top}y(t)+2x(t)^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}QRQ}x(t)\,\mathrm{d}t+H(x(t_{i+1}))-H(x(t_{i}))\bigg{)}\times$
$\bigg{(}\!\int_{t_{i}}^{t_{i+1}}\!\!\!\!\!\!u(t)B^{\top}Q\varphi(t)\!+\!2x(t)^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}QRQ}\varphi(t))\,\mathrm{d}t\!+\!H^{\prime}(x(t_{i+1}))\varphi(t_{i+1})\!-\!H^{\prime}(x(t_{i}))\varphi(t_{i})\bigg{)},$
@@
$\langle\mathcal{A}_{h}^{-}\lambda,\varphi_{h}\rangle_{X_{h}^{*},X_{h}}=\langle I^{\prime}(x_{h}),\varphi_{h}\rangle_{X_{h}^{*},X_{h}}\qquad\forall\varphi_{h}\in X_{h},$ (4.2)
@@
$\left(\begin{smallmatrix}I-h_{1}A^{\top}&-I&0&0&\cdots&0\\ 0&I-h_{2}A^{\top}&-I&0&\cdots&0\\ 0&0&I-h_{3}A^{\top}&-I&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\vdots&\vdots&0&I-h_{M}A^{\top}&-I\\ 0&0&0&0&0&I\\ \end{smallmatrix}\right)\begin{pmatrix}\lambda_{0}\\ \lambda_{1}\\ \vdots\\ \lambda_{M}\end{pmatrix}=\left(\begin{smallmatrix}\langle I^{\prime}_{\mathrm{loc}}(x),\varphi_{1}\rangle_{X_{h}^{*},X_{h}}\\ \langle I^{\prime}_{\mathrm{loc}}(x),\varphi_{2}\rangle_{X_{h}^{*},X_{h}}\\[1.99997pt] \vdots\\[6.99997pt] \langle I^{\prime}_{\mathrm{loc}}(x),\varphi_{M}\rangle_{X_{h}^{*},X_{h}}\end{smallmatrix}\right).$ (4.3)
@@
$\langle I^{\prime}_{\mathrm{loc}}(x),\varphi_{h}\rangle_{X_{h}^{*},X_{h}}$ $\approx{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}\sum_{i=1}^{M-1}\bigg{(}h_{i}(-u_{i}^{\top}y_{i}+x_{i}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}QRQ}x_{i})+H(x_{i+1})-H(x_{i})\bigg{)}\times$
$(h_{i}(u_{i}B^{\top}Q\varphi_{i}+2x_{i}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}QRQ}\varphi_{i})+H^{\prime}(x_{i+1})\varphi_{i+1}-H^{\prime}(x_{i})\varphi_{i}).$
@@
$(I_{h}^{1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda_{h}}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda_{h}})(t_{i})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda_{h}(t_{i+1})-\lambda_{h}(t_{i})}.$
@@
$=\sum_{i=1}^{M}\frac{h_{i}}{2}\langle-Ax_{h}(t_{i}),\lambda_{h}(t_{i})-\lambda_{h}(t_{i-1})\rangle+\langle x_{h}(t_{i})-x_{h}(t_{i-1}),\lambda_{h}(t_{i})-\lambda_{h}(t_{i-1})\rangle$
$\qquad+\frac{h_{i}}{2}\left(\lambda(t_{i-1})^{\top}Bu(t_{i-1})-\lambda(t_{i})^{\top}Bu(t_{i})\right).$
@@
$\eta_{i}$ $:=\Big{|}\frac{h_{i}}{2}\langle-Ax_{h}(t_{i}),\lambda_{h}(t_{i})-\lambda_{h}(t_{i-1})\rangle+\langle x_{h}(t_{i})-x_{h}(t_{i-1}),\lambda_{h}(t_{i})-\lambda_{h}(t_{i-1})\rangle$
$\qquad+\frac{h_{i}}{2}\left(\lambda(t_{i-1})^{\top}Bu(t_{i-1})-\lambda(t_{i})^{\top}Bu(t_{i})\right)\Big{|}.$ (4.4)
@@
There, we observe that the main additional computational effort compared to usual time stepping is the computation of the adjoint state. Thus, in the following section, we provide a parallelizable approximation for the backward equation (4.3).